OpenStudy (anonymous):

Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros.

OpenStudy (anonymous):

anyone plz

OpenStudy (anonymous):

come thru please

OpenStudy (campbell_st):

well what about a = 1, b = 0 and c = 0

OpenStudy (anonymous):

can this work

OpenStudy (anonymous):

2x^2+6x-5

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well it can work but the parabola is positive definite, meaning it doesn't have x intercepts.. you could try y = x^2 -6x + 5 that works

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

how can i find out the y intercept for x^2 - 6x +5

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (anonymous):

@FlamePlaysMC

OpenStudy (taramgrant0543664):

To find yintercept you set x as 0

OpenStudy (taramgrant0543664):

y=(0)^2-6(0)+5 y=?

OpenStudy (campbell_st):

to find the y-intercept let x = 0 so y = 0 - 0 + 5 so the y-intercept is y = 5 x intercepts let y = 0 so solve \[x^2 - 6x + 5 = 0\] find the factors of 5 that add to -6 both are negative then the factored form of the equation is (x - factor 1)(x - factor 2) = 0 then find the values of x that make each factor equal zero x - factor 1 = 0 x - factor 2 = 0 these values will be the x intercepts

OpenStudy (anonymous):

Ok so now that I have the y intercept i have to find out the zeros

OpenStudy (anonymous):

How would i do that

OpenStudy (anonymous):

@campbell_st